OpenStudy (anonymous):

plz help: express as a single fraction the exact value of sin 75degrees

OpenStudy (anonymous):

hint: 75 = 30 + 45

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

sum forula for sin.

OpenStudy (anonymous):

"formula"

OpenStudy (anonymous):

rely?

OpenStudy (anonymous):

idk this formula

OpenStudy (anonymous):

sin(α + β) = sin(α)cos(β) + cos(α)sin(β) http://www.purplemath.com/modules/idents.htm

OpenStudy (mathmale):

What pgpilot is referring to is the formula sin (a-b) = sin a cos b - cos a sin b. pgpilot correctly split up 75 degrees into 45 and 30 by subtraction.

OpenStudy (anonymous):

addition dude

OpenStudy (anonymous):

oh but how do u no to use tht formula specially?

OpenStudy (mathmale):

sin(α + β) = sin(α)cos(β) + cos(α)sin(β) is related... and is more appropriate .... as pgpilot has reminded this "dude." \ 75 = 45 + 30. @lyubas: stop and think for a moment. What's the next step, now that you ahve the proper formula?

OpenStudy (anonymous):

@mathmale , no offense intended!!!

OpenStudy (anonymous):

but how do u no to use tht formula specially?

OpenStudy (mathmale):

All I can say is that it comes with practice. sin 75 is hard to evaluate, but if you break up 75 into the angles 45 and 30 degrees, whose sines and cosines are easily remembered, you have a formula by which to find sin 75. sin 75 = sin ( ? + ? ) @lyubas: Can you fill in the missing quantities? @pgpilot326: Absolutely no offense taken. Thank you for putting me back on track and forgive me for jumping into your discussion with @lyubas.

OpenStudy (anonymous):

sin(75) = sin(30+45)= fill in the rest using the formula

OpenStudy (mathmale):

@lyubas: I have a job for you (that was one of my father's favorite sayings): find sin 30, sin 45, cos 30 and cos 45. You'll need them to complete this problem.

OpenStudy (anonymous):

@mathmale you're welcome to jump in anytime. you gave me quite the chuckle! @lyubas is this (the math) making sense?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

so I got sin75=45+30=sin45cos30+cos30sin45

OpenStudy (anonymous):

sin30=radical1/2 sin45 radical 2/2 cos30radical3/2

OpenStudy (anonymous):

cos45=radical2/2

OpenStudy (mathmale):

sorry to be picking on you, but some of sin75=45+30=sin45cos30+cos30sin45 is out of order. sin 75 = sin (45 + 30) = sin 45 cos 30 + (you finish)

OpenStudy (mathmale):

\[\cos 45\deg=\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (mathmale):

Right! what about sin 45 deg? cos 30 deg? cos 45 deg?

OpenStudy (anonymous):

sin30=radical1/2 sin45 radical 2/2 cos30radical3/2

OpenStudy (anonymous):

sin 75 = sin (45 + 30) = sin 45 cos 30 +cos45sin30

OpenStudy (anonymous):

correct?

OpenStudy (mathmale):

Nevermind. I didn't see that you'd already typed in your answers. Plese proceed....evaluate sin 75 deg using the formula which our friend @pgpilot326 has provided.

OpenStudy (anonymous):

yes, good job

OpenStudy (mathmale):

Yes, now you have that correct. Can you now evaluate sin 75? sin 75 = ( )*( ) + ( )*( )

OpenStudy (anonymous):

\[\sin 75=\sin 75=\sqrt{2/2}(\sqrt{3/2)}+\sqrt{2/2 (\sqrt{1/2}}\]

OpenStudy (anonymous):

i believe you have the right idea but perhaps need a little work on using latex. is this what you mean...\[\sin75=\frac{ \sqrt{2} }{ 2 }\cdot\frac{ \sqrt{3} }{ 2 }+\frac{ \sqrt{2} }{ 2 }\cdot\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

yesss lol sry

OpenStudy (anonymous):

okay, then all you have left to do is to simplify the expression on the right.

OpenStudy (anonymous):

how do I multiply the radicals?

OpenStudy (mathmale):

So, lyubas, can you do the multiplication indicated, and then, as pgpilot says, simplify the resulting expression on the right?

OpenStudy (anonymous):

\[\sqrt{a}\cdot \sqrt{b}=\sqrt{a\cdot b}\]

OpenStudy (anonymous):

or you can factor... it doesn't matter it should be the same number

OpenStudy (anonymous):

thanks for the gumball, Mickey!

OpenStudy (anonymous):

i stil dnt gt how to multiply radical2/2 amd radical3/2?

OpenStudy (anonymous):

if you notice, both terms will have a denominator of 4. \(\sqrt{2}\cdot\sqrt{3}=\sqrt{2\cdot 3}=\sqrt{6}\)

OpenStudy (anonymous):

so you should get\[\sin 75=\frac{ \sqrt{6}+\sqrt{2} }{ 4 }\]

OpenStudy (anonymous):

oh gt it radical6/4?

OpenStudy (anonymous):

sin75=raidcal8/4?

OpenStudy (anonymous):

thts what i got

OpenStudy (anonymous):

can i divide on both sides now?

OpenStudy (anonymous):

and the \(\sqrt{2}\) comes from the second term (which also has a denominator of 4). you can't add \(\sqrt{6}+sqrt{2}\) and get \(\sqrt{8}\). doesn't work that way. if it did, then \(\sqrt{4}+\sqrt{4}=\sqrt{8}\) but \(\sqrt{4}+\sqrt{4}=2=2=4\ne\sqrt{8}\)

OpenStudy (anonymous):

oops... \(\sqrt{4}+\sqrt{4}=\sqrt{8}\) but \(\sqrt{4}+\sqrt{4}=2+2=4\ne\sqrt{8}\)

OpenStudy (anonymous):

ok so what cani do next?

OpenStudy (anonymous):

now i have: sin75=radical8/4 what do i do next?

OpenStudy (anonymous):

that's it... \[\sin75=\frac{ \sqrt{6}+\sqrt{2} }{4 }\]

OpenStudy (anonymous):

ohh i gt it lol u cnt relty do anything afta so u jst leave it!

OpenStudy (anonymous):

again, no \(\sqrt{8}\) you can't combine like that. they're not like terms. here's a link to a page to refresh your memory on radicals... http://www.purplemath.com/modules/radicals3.htm

OpenStudy (anonymous):

ye ik tht was a mistake

OpenStudy (anonymous):

canu help me with another rques?

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

it's okay, but you may want to refresh your memory anyways! that's why I'm on this site actually... to stay fresh and sharp on my math skills! I'm actually leaving so let's call back @mathmale and see if that "dude" can give you a hand!

OpenStudy (anonymous):

good luck and happy learning!

OpenStudy (anonymous):

oh ok thnks for helping ur awesomew

OpenStudy (anonymous):

thanks!